OpenStudy (anonymous):

The circumference of a circle is 24 meters. Estimate it's diameter.

OpenStudy (johnweldon1993):

Well if the circumference of a circle is \(\large 2 \pi r\) We can take the fact that Circumference = 24 \[\large \frac{24}{2 \pi} = r\] And once you get that...you know that Diameter is just the radius times 2...